# Towards Faithful and Robust LLM Specialists for Evidence-Based Question-Answering

### v3

@@ Abstract
Advances towards more faithful and traceable answers of Large Language Models (LLMs) are crucial for various research and practical endeavors. One avenue in reaching this goal is basing the answers on reliable sources. However, this Evidence-Based QA has proven to work insufficiently with LLMs in terms of citing the correct sources (source quality) and truthfully representing the information within sources (answer attributability). In this work, we systematically investigate how to robustly fine-tune LLMs for better source quality and answer attributability. Specifically, we introduce a data generation pipeline with automated data quality filters, which can synthesize diversified high-quality training and testing data at scale. We further introduce four test sets to benchmark the robustness of fine-tuned specialist models. Extensive evaluation shows that fine-tuning on synthetic data improves performance on both in- and out-of-distribution. Furthermore, we show that data quality, which can be drastically improved by proposed quality filters, matters more than quantity in improving Evidence-Based QA.111 We will open-source all our codes, LLM generations, and human annotations.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023) have become the center of many cutting-edge applications due to their generalisability and information processing abilities. A typical application of LLMs is Evidence-Based Question Answering (QA), where LLMs are expected to answer questions based on provided sources and cite the sources accurately (e.g., Ni et al., 2023; Vaghefi et al., 2023; Cui et al., 2023; Liu et al., 2023a). By providing these additional sources, multiple shortcomings of standalone LLMs, such as hallucination (Ji et al., 2023) and limited knowledge capacity (Hu et al., 2023), can be addressed, thereby enhancing answer traceability (Gao et al., 2024).
@@
To address C1, we propose a data generation pipeline that synthesizes SynSciQA (Synthetic Scientific Question Answering), a well-diversified synthetic dataset for Evidence-Based QA. We further extend the pipeline with two novel quality filters to sift out low-quality synthetic data points, leading to SynSciQA+ and SynSciQA++ (see the left half of Figure 1). To address C2, we first collect an in-domain test set SynSciQA with the data generation pipeline, which shares the data distribution with the training data (i.e., SynSciQA) but covers different topics. We further collect three test sets with different distances to the training data distribution to study the OOD performance (see the right half of Figure 1).
@@ 3 Training Data Generation
Manually annotating Evidence-Based QA data that fulfills all quality dimensions is costly and lacks scalability. In this section, we introduce a novel data generation pipeline to obtain high-quality synthetic data. First, we use OpenAI LLMs to create a diverse and broad base data set of task-specific instruction-answer pairs (SynSciQA). Second, we use our quality dimensions to create data sets of higher quality (SynSciQA+ and SynSciQA++), enabling explorations on the importance of data quality (Zhou et al., 2023).
@@ 4 Evaluation Datasets
To assess the validity of the resulting models and their in- and out-of-distribution generalisability, we create a series of four evaluation datasets. The key differences are whether the data stems from a synthetic or real-world use case and how close the underlying data distribution is to the SynSciQA trainset. Our first evaluation benchmark is SynSciQA which comprises 539 samples.
Our second evaluation benchmark is GenSearch. This is a dataset adapted from Liu et al. (2023b). In this project, the authors create a dataset from posed questions to generative search engines and mark the question-relevant text part in the given source. We take this dataset and hand-evaluate 600 question-source pairs to distill full-text questions and clear corresponding sources that contain distinct variations of the information. This results in 276 question-source pairs or 106 questions with an average of 2.6 relevant sources. After retaining this dataset, we can follow the creation process of SynSciQA (see Step 4 in section 3.1). See Appendix E for more details.
We further create ChatReport. ChatReport is an open-source RAG tool that analyses companies’ (sustainability) reports (Ni et al., 2023). It uses eleven sustainability-related questions to analyze the company’s disclosure. Inherently, RAG systems’ answers rely on source paragraphs from the underlying document, i.e. the company’s report. Thus, we use the top-10 most relevant paragraphs (retrieved by ChatReport source code) as input for our system and create 110 instructions. This means we leave the structure of relevant / irrelevant sources and adopt a genuine RAG setting.
Finally, we use another RAG tool to create ClimateQA. ClimateQA333https://huggingface.co/spaces/Ekimetrics/climate-question-answering is a RAG system that answers questions based on IPCC and IPBES reports. We pose 261 climate-related questions from Welch (2022) to the system and store the outputted sources. Again, we use this data as input for our instruction form.
Figure 3 illustrates the distance between proposed test sets and SynSciQA in dimensions of use case and data distribution. ChatReport and ClimateQA are directly extracted from real applications while GenSearch is also from real research engine retrieval but with manual parsing (semi-synthetic). They also have different distribution distances to SynSciQA. GenSearch contains vastly diversified, non-scientific questions; the source texts of ChatReport contain formatting noise from sustainability reports; and ClimateQA contains nested citations in its source texts, which may influence the models’ citation correctness. Further explorations for each test set are showcased in Appendix F.
@@ 5.1 Zero-Shot Performance
| SynSciQA | GenSearch | ChatReport | ClimateQA | |||||
| Source. | Attr. | Source. | Attr. | Source. | Attr. | Source. | Attr. | |
| Llama-2-13b-chat | - | - | ||||||
| Zephyr-7b- | - | - | ||||||
| GPT-3.5 | - | - | ||||||
| GPT-4 | - | - | ||||||
@@
All models achieve lower attributability scores on non-synthetic test sets, indicating that these more realistic settings are more challenging and current LLMs are far from faithful in Evidence-Based QA. The source quality scores on GenSearch are relatively high since its questions and corresponding sources are extremely diversified (see Appendix E). Thus, it is easier to tell whether a source is relevant to a question or not.
@@ 5.2 SynSciQA Fine-Tuning
To study RQ1, we fine-tune open-sourced LLMs on SynSciQA datasets of different qualities. To control quantity when comparing quality, we randomly sample subsets of SynSciQA and SynSciQA+, leading to SynSciQA and SynSciQA+ with the same data quantity as SynSciQA++. To study RQ2, we evaluate all fine-tuned checkpoints on test sets of different distributions to see if synthetic data fine-tuning leads to overall improvement. We further calculate the correlation between in-domain (SynSciQA) and OOD (other three test sets) performance to check if in-domain performance can indicate real-world performance. All fine-tuning lasts 5 epochs and we report the performance of all epochs for two reasons: (1) we suspect that epoch number is an essential hyperparameter for OOD performance, as too many epochs may lead to overfitting to synthetic data; and (2) little previous work explores the influence of epoch number and potential overfitting in instructing tuning.
RQ1: Quality matters more than quantity. We first compare the fine-tuning performance with data of different quality, having the quantity controlled. Figure 4 shows that fine-tuning data with better source quality leads to higher source quality scores (SynSciQA+ and SynSciQA++ outperform SynSciQA). Figure 5 shows that higher data quality also leads to better attributability, where in most cases (75%) SynSciQA++ > SynSciQA+ > SynSciQA. Fine-tuning on the highest quality data even leads to comparable or better performance than GPT-4 on SynSciQA and GenSearch, and GPT-3.5-comparable performance on ChatReport and ClimateQA.
@@
RQ2.2: Synthetic data as validation set for OOD performance. We observe a fluctuating performance corresponding to fine-tuning epochs. Therefore, it is important to conduct checkpoint selection over epochs with a validation (or development) set during fine-tuning. However, performance on SynSciQA is much higher than that of other OOD test sets. Therefore, the performance on a synthetic dataset cannot directly reflect the OOD or real-world performance. But can in-domain synthetic data still be an effective development set indicating which epoch may perform best on OOD data? To answer this question, we compute the Pearson’s Correlation between performance scores of different test sets. Results are presented in Table 3, illustrating that the performance on synthetic data has a strong correlation with OOD performance. However, the correlation becomes weaker when the distribution is more distant (ClimateQA and ChatReport has a weaker correlation than GenSearch). Therefore, we conclude that synthetic data can provide a valid development set for OOD performance.
@@
| SynSciQA | & | GenSearch Src. | 0.99 | 0.97 |
| SynSciQA | & | GenSearch Attr. | 0.96 | 0.98 |
| SynSciQA | & | ChatReport Attr. | 0.94 | 0.93 |
| SynSciQA | & | ClimateQA Attr. | 0.91 | 0.94 |
RQ2.3: Overfitting does exist. We suspect that fine-tuning too many epochs may cause overfitting and reduce generalisability. So we compute the Pearson Correlation between the performance scores and epoch numbers of all settings, where positive correlations indicate benefit from more epochs and negative correlations indicate overfits. Results are visualized in Figure 8, showing an overfitting trend for the majority of settings. Therefore, fine-tuning too many epochs may lead to a suboptimal performance. But we do not observe the fine-tuning overfits more to SynSciQA than others. We attribute this to SynSciQA containing different scientific topics from the training data.
@@ 7 Conclusion
Data Privacy or Bias: There are no data privacy issues or bias against certain demographics with regard to the data collected from real-world applications and LLM generations. All artifacts we use are under creative common licence. We also notice no ethical risks associate with this work
@@ Appendix A The Challenge of Helpfulness
| SynSciQA | GenSearch | ChatReport | ClimateQA | |||||||||||||||||||||||||||||||||
@@
Collectively, we view the exact measurement of helpfulness as a challenge for future work. But we argue that our two employed evaluation metrics already address helpfulness in Evidence-Based QA to a satisfactory degree.
@@
Answer by GPT-4 |
None of the provided sources address the interaction between natural selection and genetic drift. |
Answer by Llama-13b-chat |
Sure, I’d be happy to help! Natural selection and genetic drift are both important mechanisms that drive the evolution of species. While natural selection is the process by which individuals with advantageous traits are more likely to survive and reproduce, genetic drift is the random change in the frequency of alleles in a population over time (Morse et al., 2018, p. 67). The interaction between natural selection and genetic drift can be complex, as both processes can influence the evolution of a species in different ways. For example, natural selection can lead to the fixation of advantageous alleles in a population, while genetic drift can result in the loss or fixation of neutral or deleterious alleles (Hudsonsonian Institution, 2017, p. 28). Additionally, […] |
Answer by Llama-13b-chat fine-tuned for two epochs on SynSciQA++ |
None of the provided sources address the question "How does natural selection interact with genetic drift?" Therefore, an answer cannot be given based on these sources. |
@@
Answer by GPT-4 |
The current top-ranked female tennis player in the world is Iga Swiatek (Online2602022, 2019, p.8). |
Answer by Zephyr-7b- |
As of the sources provided, Iga Swiatek currently holds the top ranking in women’s singles tennis. This information can be found in the source "Online2602022" (Williams, 2022, p.8). Ashleigh Barty, the previous world No.1, retired early in 2022, which led to Swiatek’s ascension to the top spot. (Williams, 2022, p.8) |
Answer by Zephyr-7b- fine-tuned for two epochs on SynSciQA++ |
The top ranked female tennis player in the world is Iga Swiatek (Online2602022, 2019, p.8). |
@@ Appendix B Attributability Score Details
We use SpaCy (Honnibal and Montani, 2017) to split answers into sentences. Unattributable answer sentences caused by missing citations or wrong citation format can be easily identified through golden heuristics, for example, matching citations with actual source names. However, it is hard for a heuristic-based method to judge whether a statement is entailed by the cited source. Previous work proposes to use NLI models to predict entailment (Honovich et al., 2022; Gao et al., 2023; Yue et al., 2023). Among them, Yue et al. (2023) aggregate the largest NLI training set and conduct extensive analyses to explore the best practice of attributability prediction. Therefore, we rely on their results to select models for the attributability score. The two best-performing checkpoints are Flan-t5-XL and Flan-t5-XXL 444https://huggingface.co/osunlp/attrscore-flan-t5-xxl. When inferencing with these checkpoints, we follow the prompt template in Yue et al. (2023) and use greedy generation. We aggregate the prediction of both models to improve the precision since false positives are more harmful than false negatives in the task of judging LLM faithfulness.

### v4

@@ Abstract
Advances towards more faithful and traceable answers of Large Language Models (LLMs) are crucial for various research and practical endeavors. One avenue in reaching this goal is basing the answers on reliable sources. However, this Evidence-Based QA has proven to work insufficiently with LLMs in terms of citing the correct sources (source quality) and truthfully representing the information within sources (answer attributability). In this work, we systematically investigate how to robustly fine-tune LLMs for better source quality and answer attributability. Specifically, we introduce a data generation pipeline with automated data quality filters, which can synthesize diversified high-quality training and testing data at scale. We further introduce four test sets to benchmark the robustness of fine-tuned specialist models. Extensive evaluation shows that fine-tuning on synthetic data improves performance on both in- and out-of-distribution. Furthermore, we show that data quality, which can be drastically improved by proposed quality filters, matters more than quantity in improving Evidence-Based QA.111 All our codes, LLM generations, and human annotations are accessible through https://github.com/EdisonNi-hku/Robust_Evidence_Based_QA.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023) have become the center of many cutting-edge applications due to their generalisability and information processing abilities. A typical application of LLMs is in Evidence-Based Question Answering (QA), where LLMs are expected to answer questions based on provided sources and cite the sources accurately (e.g., Ni et al., 2023; Vaghefi et al., 2023; Cui et al., 2023; Liu et al., 2023a). By providing these additional sources, multiple shortcomings of standalone LLMs, such as hallucination (Ji et al., 2023) and limited knowledge capacity (Hu et al., 2023), can be addressed, thereby enhancing answer traceability (Gao et al., 2024).
@@
To address C1, we propose a data generation pipeline that synthesizes SynSciQA (Synthetic Scientific Question Answering), a well-diversified synthetic dataset for Evidence-Based QA, following prior work on data distillation for instruction tuning (e.g., Honovich et al., 2023; Tunstall et al., 2023). We further extend the pipeline with two novel quality filters to sift out low-quality synthetic data points, leading to SynSciQA+ and SynSciQA++ (see the left half of Figure 1). To address C2, we first collect an in-domain test set SynSciQAtest with the data generation pipeline, which shares the data distribution with the training data (i.e., SynSciQA) but covers different topics. We further collect three test sets with different distances to the training data distribution to study the OOD performance (see the right half of Figure 1).
@@ 3 Training Data Generation
Manually annotating Evidence-Based QA data that fulfills all quality dimensions is costly and lacks scalability. In this section, we introduce a novel data generation pipeline to obtain high-quality synthetic data. First, we use OpenAI LLMs to create a diverse and broad base data set of task-specific instruction-answer pairs (SynSciQA) following the structural approach of prior work for data distillation (e.g., Honovich et al., 2023; Tunstall et al., 2023). Second, we use our quality dimensions to create data sets of higher quality (SynSciQA+ and SynSciQA++), enabling explorations on the importance of data quality (Zhou et al., 2023).
@@ 4 Evaluation Datasets
To assess the validity of the resulting models and their in- and out-of-distribution generalisability, we create a series of four evaluation datasets. The key differences are whether the data stems from a synthetic or real-world use case and how close the underlying data distribution is to the SynSciQA trainset. Our first evaluation benchmark is SynSciQAtest which comprises 539 samples.
Our second evaluation benchmark is GenSearchtest. This is a dataset adapted from Liu et al. (2023b). In this project, the authors create a dataset from posed questions to generative search engines and mark the question-relevant text part in the given source. We take this dataset and hand-evaluate 600 question-source pairs to distill full-text questions and clear corresponding sources that contain distinct variations of the information. This results in 276 question-source pairs or 106 questions with an average of 2.6 relevant sources. After retaining this dataset, we can follow the creation process of SynSciQA (see Step 4 in section 3.1). See Appendix E for more details.
We further create ChatReporttest. ChatReport is an open-source RAG tool that analyses companies’ (sustainability) reports (Ni et al., 2023). It uses eleven sustainability-related questions to analyze the company’s disclosure. Inherently, RAG systems’ answers rely on source paragraphs from the underlying document, i.e. the company’s report. Thus, we use the top-10 most relevant paragraphs (retrieved by ChatReport source code) as input for our system and create 110 instructions. This means we leave the structure of relevant / irrelevant sources and adopt a genuine RAG setting.
Finally, we use another RAG tool to create ClimateQAtest. ClimateQA333https://huggingface.co/spaces/Ekimetrics/climate-question-answering is a RAG system that answers questions based on IPCC and IPBES reports. We pose 261 climate-related questions from Welch (2022) to the system and store the outputted sources. Again, we use this data as input for our instruction form.
Figure 3 illustrates the distance between proposed test sets and SynSciQAtest in dimensions of use case and data distribution. ChatReporttest and ClimateQAtest are directly extracted from real applications while GenSearchtest is also from real research engine retrieval but with manual parsing (semi-synthetic). They also have different distribution distances to SynSciQA. GenSearchtest contains vastly diversified, non-scientific questions; the source texts of ChatReporttest contain formatting noise from sustainability reports; and ClimateQAtest contains nested citations in its source texts, which may influence the models’ citation correctness. Further explorations for each test set are showcased in Appendix F.
@@ 5.1 Zero-Shot Performance
| SynSciQAtest | GenSearchtest | ChatReporttest | ClimateQAtest | ||||||||
| Source. | Attr. | Source. | Attr. | Source. | Attr. | Source. | Attr. | ||||
| Llama-2-13b-chat | - | - | |||||||||
| Zephyr-7b- | - | - | |||||||||
| GPT-3.5 | - | - | |||||||||
| GPT-4 | - | - | |||||||||
@@
All models achieve lower attributability scores on non-synthetic test sets, indicating that these more realistic settings are more challenging and current LLMs are far from faithful in Evidence-Based QA. The source quality scores on GenSearchtest are relatively high since its questions and corresponding sources are extremely diversified (see Appendix E). Thus, it is easier to tell whether a source is relevant to a question or not.
@@ 5.2 SynSciQA Fine-Tuning
To study RQ1, we fine-tune open-sourced LLMs on SynSciQA datasets of different qualities. To control quantity when comparing quality, we randomly sample subsets of SynSciQA and SynSciQA+, leading to SynSciQAS and SynSciQA+S with the same data quantity as SynSciQA++. To study RQ2, we evaluate all fine-tuned checkpoints on test sets of different distributions to see if synthetic data fine-tuning leads to overall improvement. We further calculate the correlation between in-domain (SynSciQAtest) and OOD (other three test sets) performance to check if in-domain performance can indicate real-world performance. All fine-tuning lasts 5 epochs and we report the performance of all epochs for two reasons: (1) we suspect that epoch number is an essential hyperparameter for OOD performance, as too many epochs may lead to overfitting to synthetic data; and (2) little previous work explores the influence of epoch number and potential overfitting in instructing tuning.
RQ1: Quality matters more than quantity. We first compare the fine-tuning performance with data of different quality, having the quantity controlled. Figure 4 shows that fine-tuning data with better source quality leads to higher source quality scores (SynSciQA+S and SynSciQA++ outperform SynSciQAS). Figure 5 shows that higher data quality also leads to better attributability, where in most cases (75%) SynSciQA++ > SynSciQA+S > SynSciQAS. Fine-tuning on the highest quality data even leads to comparable or better performance than GPT-4 on SynSciQAtest and GenSearchtest, and GPT-3.5-comparable performance on ChatReporttest and ClimateQAtest.
@@
RQ2.2: Synthetic data as validation set for OOD performance. We observe a fluctuating performance corresponding to fine-tuning epochs. Therefore, it is important to conduct checkpoint selection over epochs with a validation (or development) set during fine-tuning. However, performance on SynSciQAtest is much higher than that of other OOD test sets. Therefore, the performance on a synthetic dataset cannot directly reflect the OOD or real-world performance. But can in-domain synthetic data still be an effective development set indicating which epoch may perform best on OOD data? To answer this question, we compute the Pearson’s Correlation between performance scores of different test sets. Results are presented in Table 3, illustrating that the performance on synthetic data has a strong correlation with OOD performance. However, the correlation becomes weaker when the distribution is more distant (ClimateQA and ChatReport has a weaker correlation than GenSearch). Therefore, we conclude that synthetic data can provide a valid development set for OOD performance.
@@
| SynSciQAtest | & | GenSearchtest Src. | 0.99∗∗ | 0.97∗∗ |
| SynSciQAtest | & | GenSearchtest Attr. | 0.96∗∗ | 0.98∗∗ |
| SynSciQAtest | & | ChatReporttest Attr. | 0.94∗∗ | 0.93∗∗ |
| SynSciQAtest | & | ClimateQAtest Attr. | 0.91∗∗ | 0.94∗∗ |
RQ2.3: Overfitting does exist. We suspect that fine-tuning too many epochs may cause overfitting and reduce generalisability. So we compute the Pearson Correlation between the performance scores and epoch numbers of all settings, where positive correlations indicate benefit from more epochs and negative correlations indicate overfits. Results are visualized in Figure 8, showing an overfitting trend for the majority of settings. Therefore, fine-tuning too many epochs may lead to a suboptimal performance. But we do not observe the fine-tuning overfits more to SynSciQA than others. We attribute this to SynSciQAtest containing different scientific topics from the training data.
@@ 6 Related Work
7 Discussions and Future Work
Broader Impact: The aim of this work is to build a basis for constantly improving open-source LLMs in Evidence-Based QA, which is important for the practical community where RAG is heavily employed in applications. The NLP research community may also find our work inspiring in mitigating LLM hallucination: our proposed paradigm for Evidence-Based QA requires all answer sentences to be grounded by in-context sources. Such controlled generation makes hallucination detection much easier leveraging entailment models. Human evaluations in Appendix D and Appendix J also prove the potential of NLI-based hallucination detection. Future Work: For research, we will continue improving open-sourced LLM’s performance on Evidence-Based QA: For example, (1) continuing fine-tuning existing instruction-fine-tuned checkpoints on RLHF alignment stages; (2) generalizing LLM specialists to other templates to study the trade-off between specialization and generalization; and (3) exploring how to leverage LLM parametric knowledge with attributability. For the practical community, we will continuously benchmark new LLMs on our datasets. At the same time, we aim to make the resulting models accessible for the practical community444See the updates on https://github.com/EdisonNi-hku/Robust_Evidence_Based_QA.. Furthermore, we outline that the training data for this project was mainly (75%) distilled from GPT-3.5 instead of GPT-4, making it more accessible for low-budget RAG development. More powerful generic LLMs for data distillation may improve the results even more.
@@ 8 Conclusion
Data Privacy or Bias: There are no data privacy issues or bias against certain demographics with regard to the data collected from real-world applications and LLM generations. All artifacts we use are under a creative common license. We also notice no ethical risks associated with this work.
@@ Appendix A The Challenge of Helpfulness
| SynSciQAtest | GenSearchtest | ChatReporttest | ClimateQAtest | |||||||||||||||||||||||||||||||||
@@
Collectively, we view the exact measurement of helpfulness as a challenge for future work. However, we argue that our two employed evaluation metrics already address helpfulness in Evidence-Based QA to a satisfactory degree.
To improve the helpfulness evaluation, future work could try to identify dimensions of helpfulness that are perpendicular to source quality and answer attributability and evaluate them with the help of LLMs. One dimension could be friendliness or the degree how well the question is addressed. However, as outlined, these dimensions might stand in conflict with the two quality dimensions introduced in this work. Thus, investigating these trade-offs could present an interesting new direction.
@@
| Answer by GPT-4 | None of the provided sources address the interaction between natural selection and genetic drift. |
| Answer by Llama-13b-chat | Sure, I’d be happy to help! Natural selection and genetic drift are both important mechanisms that drive the evolution of species. While natural selection is the process by which individuals with advantageous traits are more likely to survive and reproduce, genetic drift is the random change in the frequency of alleles in a population over time (Morse et al., 2018, p. 67). The interaction between natural selection and genetic drift can be complex, as both processes can influence the evolution of a species in different ways. For example, natural selection can lead to the fixation of advantageous alleles in a population, while genetic drift can result in the loss or fixation of neutral or deleterious alleles (Hudsonsonian Institution, 2017, p. 28). Additionally, […] |
| Answer by Llama-13b-chat fine-tuned for two epochs on SynSciQA++ | None of the provided sources address the question "How does natural selection interact with genetic drift?" Therefore, an answer cannot be given based on these sources. |
@@
| Answer by GPT-4 | The current top-ranked female tennis player in the world is Iga Swiatek (Online2602022, 2019, p.8). |
| Answer by Zephyr-7b- | As of the sources provided, Iga Swiatek currently holds the top ranking in women’s singles tennis. This information can be found in the source "Online2602022" (Williams, 2022, p.8). Ashleigh Barty, the previous world No.1, retired early in 2022, which led to Swiatek’s ascension to the top spot. (Williams, 2022, p.8) |
| Answer by Zephyr-7b- fine-tuned for two epochs on SynSciQA++ | The top ranked female tennis player in the world is Iga Swiatek (Online2602022, 2019, p.8). |
@@ Appendix B Attributability Score Details
We use SpaCy (Honnibal and Montani, 2017) to split answers into sentences. Unattributable answer sentences caused by missing citations or wrong citation format can be easily identified through golden heuristics, for example, matching citations with actual source names. However, it is hard for a heuristic-based method to judge whether a statement is entailed by the cited source. Previous work proposes to use NLI models to predict entailment (Honovich et al., 2022; Gao et al., 2023; Yue et al., 2023). Among them, Yue et al. (2023) aggregate the largest NLI training set and conduct extensive analyses to explore the best practice of attributability prediction. Therefore, we rely on their results to select models for the attributability score. The two best-performing checkpoints are Flan-t5-XL and Flan-t5-XXL 555https://huggingface.co/osunlp/attrscore-flan-t5-xxl. When inferencing with these checkpoints, we follow the prompt template in Yue et al. (2023) and use greedy generation. We aggregate the prediction of both models to improve the precision since false positives are more harmful than false negatives in the task of judging LLM faithfulness.